# Products, Abstractions and Inclusions of Causal Spaces

### v1

@@ Appendix C Equivalence of causal mechanism definitions
Let us here briefly comment on the notation and equivalence of the notions $K_{S}(\omega,A)$ to $K_{S}(\omega_{S},A)$. To clarify the equivalence let us recall the factorisation lemma.
Lemma C.1 (Factorisation Lemma).
@@
We apply this where for $T$ we use the projection $\pi_{S}:(\Omega,\mathscr{H})\to(\Omega_{S},\mathscr{H}_{S})$. Then, by the factorisation lemma, a $\mathscr{H}_{S}$-measurable map $K_{S}(\cdot,A):\Omega\to[0,1]$ gives rise to a map $K_{S}^{\prime}(\cdot,A):\Omega_{S}\to[0,1]$ such that
@@
Moreover, this maps is unique because $\pi_{S}$ is surjective. On the other hand, given $K_{S}^{\prime}(\cdot,A):\Omega_{S}\to[0,1]$, we can just define $K_{S}(\cdot,A):\Omega\to[0,1]$ by (5). For this reason $K_{S}(\omega,A)$ and $K_{S}(\omega_{S},A)$ can be used equivalently when identified as explained here and with a slight abuse of notation we will use both depending on the context without indicating the different domains.

### v2

@@ Appendix C Equivalence of causal mechanism definitions
Let us here briefly comment on the notation and equivalence of the notions $K_{S}(\omega,A)$ to $K_{S}(\omega_{S},A)$. To clarify the equivalence let us recall the factorization lemma.
Lemma C.1 (Factorization Lemma).
@@
We apply this where for $T$ we use the projection $\pi_{S}:(\Omega,\mathscr{H})\to(\Omega_{S},\mathscr{H}_{S})$. Then, by the factorization lemma, a $\mathscr{H}_{S}$ measurable map $K_{S}(\cdot,A):\Omega\to[0,1]$ gives rise to a map $K_{S}^{\prime}(\cdot,A):\Omega_{S}\to[0,1]$ such that
@@
Moreover, this maps is unique because $\pi_{S}$ is surjective. On the other hand, give $K_{S}^{\prime}(\cdot,A):\Omega_{S}\to[0,1]$ we can just define $K_{S}(\cdot,A):\Omega\to[0,1]$ by (11). For this reason $K_{S}(\omega,A)$ and $K_{S}(\omega_{S},A)$ can be used equivalently when identified as explained here and with a slight abuse of notation we will use both depending on the context without indicating the different domains.